Department E had 4,000 units in work in process that were 40% completed at the beginning of the period at a cost of $12,500. During the period, 14,000 units of direct materials were added at a cost of $28,700 and 15,000 units were completed. At the end of the period, 3,000 units were 75% completed. All materials are added t the beginning of the process. Direct labor was $32,450, and factory overhead was $18,710. The number of equivalent units of production for the period for conversion if the first-in, first-out method is used to cost inventories was

A. 14,850

B. 14,650

C. 15,650

D. 14,150
Answer #1

Unit transferred out = 4000+14000-3000 = 15000

Equivalent unit of conversion = (4000*60%)+11000+(3000*75%) = 15650

So answer is c) 15650